OpenStudy (anonymous):

find cos pi/4 without using a calculator.

OpenStudy (amistre64):

you would have to know what the right triangle for pi/4 looks like

OpenStudy (anonymous):

sqrt 2/2

OpenStudy (amistre64):

pi = 180 when converting radians to degrees pi/4 radians = 180/4 degrees = 45 degrees

OpenStudy (amistre64):

a 45 degree angled right triangle can be fashioned out of a square that is cut diagonally in half

OpenStudy (amistre64):

since a square has all sides equal; each leg of the rt trianlge is the same

OpenStudy (amistre64):

to find the diagonal we can use the pythag thrm such that: s^2 + s^2 = diag^2 2s^2 = diag^2 diag = sqrt(2s^2) = (s)sqrt(2) when the sides = 1 it is quite easy to see that the diag = sqrt(2)

OpenStudy (amistre64):

now cosine is defined as the adjacent side over the hypotenuse; lets use the easy measurement of 1 for each side then; cos(45) = 1/sqrt(2) which can be reshaped into: sqrt(2)/2 if need be for those sticklers out there who want to say that radicals dont belong in the bottom of a fraction :)

OpenStudy (anonymous):

amistre64, are you calling me a stickler? lol

OpenStudy (amistre64):

like a cactus ;)